OpenStudy (anonymous):

Rachel and Sam stopped at the donut stop before work. Rachel bought a protein smoothie and 2 boxes of donut holes for $11.65. Sam bought 2 protein smoothies and 1 box of donut holes for $10.55. Use the substitution method to solve a system of equations relating S, the cost of a smoothie, and D, the cost of a box of donut holes. S + 2D = 11.65 2S + D = 10.55 How much does each item cost? Select one: a. Smoothie $3.70; box of donut holes $3.70 b. Smoothie $4.25; box of donut holes $3.15 c. Smoothie $3.15; box of donut holes $4.25 d. Smoothie $2.95; box of donut holes $4.35

OpenStudy (anonymous):

(: shpanks

OpenStudy (dan815):

okay so sub method means we have to first isolate an equation in terms of 1 variable S+2D=11.65 S=11.65- 2D Now we sub this into the 2nd equation 2S + D = 10.55 2*(11.65-2D)+D=10.55 <---- we can simplify this and solve for D 23.3-4D+D=10.55 -3D=-12.75 D=12.75/3=4.25

OpenStudy (dan815):

S+2D=11.65 S+2(4.25)=11.65 S=11.65-8.5=3.15